A theater has 32 rows of seats. The first row has 24 seats. Each of the remaining rows increases by 2 seats. How many seats are there in the theater?
know: tn=32
a=24
r=2
find: n
So wouldn't I use the formula tn=a(r) to the exponent n-1?

Answer by scott8148(6628)

this is arithmetic (not geometric) ___ d = 2 (not r)

1st row -- 24 ___ 32nd row -- 24 + 2(31) = 86

average ___ (24 + 86) / 2 = 55

32 * 55 = 1760